At Dive Shop Inc, regular customers place about 4.0 orders per year. These orders average $93 in retail value, and contribution margin percent is typically 41%. Acquiring a new customer costs Dive Shop about $135, and Dive Shop also spends $43 per customer per year on contests, deals, and communications to keep them coming back. Dive Shop's retention rate is 83%, and the corporate discount rate is 11%. Calculate the net present value of 1 regular customer at Dive Shop. Rounding: penny. Remember, the answer may be negative.

SHOW EVERY STEP YOU DID TO FIND ANSWER.
Answer #1

Calculation of One customer:

Sales $372 (93*4)

contribution $ 152.52 (372 * 41%)

Cost of Acquiring Customer $43

Net Gain from one Customer $109.52

Retention ratio (83%) $90.9016

NPV discount factor 11 % $81.8933 (90.9016/1.11)
Answer #2

To calculate the net present value (NPV) of 1 regular customer at Dive Shop, we need to consider the cash flows associated with that customer over a certain period, and then discount these cash flows to their present value using the corporate discount rate.

Let's break down the cash flows and calculate the NPV step by step:

  1. Contribution Margin per Order: Contribution Margin per Order = Retail value * Contribution Margin percent Contribution Margin per Order = $93 * 0.41 = $38.13

  2. Contribution Margin per Year (per customer): Contribution Margin per Year = Contribution Margin per Order * Number of orders per year Contribution Margin per Year = $38.13 * 4.0 = $152.52

  3. Total Revenue per Year (per customer): Total Revenue per Year = Retail value * Number of orders per year Total Revenue per Year = $93 * 4.0 = $372

  4. Total Costs per Year (per customer): Total Costs per Year = Acquiring cost + Communication cost Total Costs per Year = $135 + $43 = $178

  5. Net Cash Flow per Year (per customer): Net Cash Flow per Year = Total Revenue per Year - Total Costs per Year Net Cash Flow per Year = $372 - $178 = $194

  6. Retained Customers (83% of the customers will be retained in the next year): Retained Customers = 0.83 * 1 (for 1 regular customer) = 0.83

  7. Cash Flow in the next year: Cash Flow in the next year = Net Cash Flow per Year * Retained Customers Cash Flow in the next year = $194 * 0.83 ≈ $160.82

  8. Calculate NPV using the formula: NPV = (Cash Flow in the next year) / (1 + Discount Rate)

Discount Rate = 11% = 0.11 (as a decimal)

NPV = $160.82 / (1 + 0.11) NPV ≈ $144.66 (rounded to the nearest penny)

The NPV of 1 regular customer at Dive Shop is approximately $144.66 (rounded to the nearest penny).

answered by: Hydra Master